OpenStudy (anonymous):

CALCULUS TROUBLE :< A manufacturer can produce digital records at a cost of 50 a piece. It is estimated that if the recorders are sold for "p" a piece, consumers will buy q=120-p recorders each month. Please express the manufacturer's profit P as a function of q.

OpenStudy (anonymous):

You want to get p on its own on one side of the equals sign. Firstly, you have to take 120 from both sides.

OpenStudy (anonymous):

p + q = 120 => p = 120 - q. p(q) = 120 - q.

OpenStudy (anonymous):

wow thanks. another question, same problem: what is the average rate of profit obtained as the level of production increases from q=0 to q=20?

OpenStudy (anonymous):

Once you have \(p(q)=120-q\), just substitute in the two values of q (i.e. find p(0), and p(20)), then take their average (add them, then divide by 2).